A wave function is given to be πœ“(π‘₯) = 𝑁( π‘₯ /π‘₯0 )^ 𝑛𝑒 βˆ’π‘₯/π‘₯0, 𝑁, 𝑛, π‘₯0 π‘Žπ‘Ÿπ‘’ π‘π‘œπ‘›π‘ π‘‘π‘Žπ‘›π‘‘π‘  From Schrodinger wave equation find potential V(x) and the energy E for which πœ“(π‘₯) is an Eigen function.


1
Expert's answer
2020-12-28T10:27:18-0500

Let's write Schrodinger equation applied to "\psi(x)" :

"-\frac{\hbar^2}{2m} \psi'' +V\psi = E\psi"

First calculate "\psi''" :

"\psi''=N e^{-x/x_0} (\frac{n(n-1)}{x_0^2} (\frac{x}{x_0})^{n-2}-2\frac{n}{x_0^2}(\frac{x}{x_0})^{n-1}+\frac{1}{x_0^2}(\frac{x}{x_0})^n)"

"\psi''=Ne^{-x/x_0} (\frac{x}{x_0})^n (\frac{1}{x_0^2}-\frac{2n}{xx_0}+\frac{n(n-1)}{x^2}) = \psi (\frac{1}{x_0^2}-\frac{2n}{xx_0}+\frac{n(n-1)}{x^2})"

Now we substitue this expression into the Schrodinger equation :

"(\frac{-\hbar^2}{2m}(\frac{1}{x_0^2}-\frac{2n}{xx_0}+\frac{n(n-1)}{x^2})+V)\psi = E\psi"

As "\psi\neq 0" everywhere except for "x=0", we can simplify by "\psi" and find

"\frac{-\hbar^2}{2m}(\frac{1}{x_0^2}-\frac{2n}{xx_0}+\frac{n(n-1)}{x^2})+V = E"

"V=E+\frac{\hbar^2}{2m}(\frac{1}{x_0^2}-\frac{2n}{xx_0}+\frac{n(n-1)}{x^2})"

As "V\to 0" when "|x|\to+\infty" , we find that

"E=-\frac{\hbar^2}{2mx_0^2}"

"V=\frac{\hbar^2}{2m}(-\frac{2n}{xx_0}+\frac{n(n-1)}{x^2})"